OpenStudy (anonymous):

Mr. McGregor is pouring concrete for the floor of a circular building. The floor of the building will have a diameter of 47 feet. What is the area of the floor of the building to the nearest square foot? Use 3.14 for p

OpenStudy (pfenn1):

Area (A) of a circle is (pi) r^2 where r = radius = D/2 where D = diameter = 47

OpenStudy (anonymous):

47(3.14) = answer feet squared

OpenStudy (pfenn1):

A=pi * (D/2)^2

OpenStudy (anonymous):

oh whoops sorry my answer is wrong.

OpenStudy (anonymous):

so what is the correct answer ???

OpenStudy (pfenn1):

A = (3.14) (47/2)^2 = (3.14) (23.5)^2